OpenStudy (anonymous):

Positive difference between x/y and x/y-1

OpenStudy (anonymous):

@LarsEighner @ikram002p

OpenStudy (ikram002p):

you mean \(\dfrac{x}{y}-\dfrac{x}{y-1}\ =? )

OpenStudy (anonymous):

heres the full problem Angelo makes x dollars for y hours sara makes the same amount while working one hour less what is the positive difference between their wages

OpenStudy (anonymous):

here

OpenStudy (ikram002p):

since sara make more so \(\dfrac{x}{y-1}-\dfrac{x}{y} >0\)

OpenStudy (ikram002p):

i think this is the correct answer :)

OpenStudy (ikram002p):

A

OpenStudy (anonymous):

It is! Thanks :)

OpenStudy (ikram002p):

np :)